OpenStudy (anonymous):

Find the perimeter of a rectangle if three of its vertices are (5,-2), (-3,-2), (-3,3)

OpenStudy (anonymous):

wht is it tht youre not understanding

OpenStudy (anonymous):

im new to this how do i find the perimeter do i find the missing point or what

OpenStudy (jdoe0001):

notice the plotted points notice their coordinates keep in mind that is a rectangle a rectangle has to OPPOSITE EQUAL sides so the bottom and top sides must be equal the left and right side must also be equal so where do you think is the last coordinate?

OpenStudy (anonymous):

(5,3)

OpenStudy (jdoe0001):

yes

OpenStudy (jdoe0001):

now just find the distances, add them all up, and that's the perimeter of that rectangle

OpenStudy (anonymous):

so i ad the x and y together (4,2)

OpenStudy (jdoe0001):

well... how many units from (-3, -2) to say (5, -2) ?

OpenStudy (jdoe0001):

if you look at the picture, you can pretty much count them

OpenStudy (anonymous):

8,

OpenStudy (jdoe0001):

so the bottom line is 8 units the OPPOSITE SIDE will also be 8 units 8 + 8 then add the other 2 sides, the left and right sides to get the perimeter :)

OpenStudy (anonymous):

the perimeter is 16 right

OpenStudy (jdoe0001):

well, the top and bottom lines are 8+8=16 but you still need to add the left and right lines too

OpenStudy (anonymous):

10 so its 16,10

OpenStudy (jdoe0001):

yeap 8+8+5+5 = 16 + 10 = 26

OpenStudy (anonymous):

ok the perimeter is 26

OpenStudy (jdoe0001):

yeap

OpenStudy (anonymous):

thanks